A woman walks at 2.1 m/s due East. After 199 seconds, she changes her direction by 20° to the South at 2.9 m/s for 187 seconds. What is the distance between where she started and where she ended?
1
Expert's answer
2020-12-14T07:17:05-0500

Find her displacement due East:


"E=v_Et_E."

Her displacement due South:


"S=v_St_S."

The distance between where she started and where she ended:


"d=\sqrt{E^2+S^2}=\sqrt{(v_Et_E)^2+(v_St_S)^2},\\
d=\sqrt{(2.1\cdot199)^2+(2.9\cdot187)^2}=685\text{ m}."